# cuPDLP+: A Further Enhanced GPU-Based First-Order Solver for Linear Programming

### v1

@@ 4 Numerical experiments
In this section, we compare the numerical performance of cuPDLP+ with cuPDLP. We first describes the setup of the experiments, followed by presents the numerical results on LP relaxations of instances from MIPLIB 2017 collection.
Benchmark datasets. We conduct numerical experiments on a widely used LP benchmark dataset: MIP Relaxations, which consists of 383 instances derived from root-node LP relaxations of MIPLIB 2017 [9]. The 383 instances in the MIP relaxations dataset are selected based on the same filtering criteria outlined in [12]. Based on the number of non-zeros of the constraint matrices, we further categorize these 383 instances into three groups, as summarized in Table 1.
@@
Software. We implement cuPDLP+ in Julia, utilizing CUDA.jl as the interface for working with NVIDIA CUDA GPUs. We compare the performance of cuPDLP+ with cuPDLP. The running time of both cuPDLP+ and cuPDLP is measured after pre-compilation in Julia.
Computing environment. We use NVIDIA H100-SXM-80GB GPU, with CUDA 12.4. The experiments are performed in Julia 1.11.5.
Initialization. cuPDLP+ uses all-zero vectors as the initial starting points.
Optimality termination criteria. cuPDLP+ terminate when the relative KKT error is no greater than the termination tolerance $\epsilon\in(0,\infty)$:
@@
The termination criteria are checked for the original LP instance, not the preconditioned ones, so that the termination is not impacted by the preconditioning. In the experiment, we set $\epsilon=10^{-4}$ for moderate accuracy and $\epsilon=10^{-8}$ for high accuracy.
Time limit. We apply a time limit of 3600 seconds to small and medium-sized problems and 18000 seconds to large-scale problems.
@@
Tables 2 and 3 compare the performance of cuPDLP and cuPDLP+ on 383 MIPLIB instances, with and without Gurobi presolve, respectively. Several key observations can be made:
  • Across all four settings, cuPDLP+ consistently outperforms cuPDLP in terms of overall numerical performance, solving more instances and achieving notable reductions in running time.
  • The performance gain of cuPDLP+ is especially significant for Small and Medium instances, while the improvement is less significant for Large instances, likely due to the high variance among the relatively small number of Large instances.
Table 4 reports the speedup of cuPDLP+ over cuPDLP. In particular, the first row focuses on hard instances, defined as those requiring at least 10 seconds to solve using either cuPDLP or cuPDLP+. In various scenarios, cuPDLP+ achieves a 2× to 4× speedup. The improvement is more substantial when Gurobi presolve is enabled compared to the case without presolve. The speedup is also more significant on harder instances. For example, in the high-accuracy setting with presolve, cuPDLP+ achieves a 2.9× speedup on all instances and a 4.63× speedup on hard instances.
@@
$\mathbf{10^{-4}}$ cuPDLP 266 8.81 91 11.55 19 77.43 376 11.07
[…]
$\mathbf{10^{-8}}$ cuPDLP 260 24.80 87 36.90 17 208.66 364 31.22
[…]
@@
$\mathbf{10^{-4}}$ cuPDLP 268 5.11 92 9.03 19 26.74 379 6.75
[…]
$\mathbf{10^{-8}}$ cuPDLP 264 18.50 90 29.40 19 63.68 373 22.42
[…]
@@
Without
Presolve
With
Presolve
1e-4 1e-8 1e-4 1e-8
Hard 3.30 3.35 4.05 4.63
Overall 2.01 2.28 2.49 2.90
Table 4: Summarization of cuPDLP+ speedup over cuPDLP. The first row reports the speedup of cuPDLP+ on instances where cuPDLP or cuPDLP+ takes more than 10 seconds to solve, and the second row summarizes the overall speedup across all instances.
@@ Acknowledgement
Haihao Lu is supported by AFOSR Grant No. FA9550-24-1-0051 and ONR Grant No. N000142412735. Zedong Peng is supported by ONR Grant No. N000142412735. Jinwen Yang is supported by AFOSR Grant No. FA9550-24-1-0051.

### v4

@@ 4 Numerical experiments
In this section, we compare the numerical performance of cuPDLPx with cuPDLP. We first describe the setup of the experiments, followed by presenting the numerical results on LP relaxations of instances from the MIPLIB 2017 collection and Mittelmann’s LP benchmark.
Benchmark datasets. We conduct numerical experiments on two widely used LP benchmark datasets: MIP Relaxations, which consists of 383 instances derived from root-node LP relaxations of MIPLIB 2017 [12], and Mittelmann’s LP benchmark dataset [19], which contains 49 LP instances. In particular, the 383 instances in the MIP relaxations dataset are selected based on the same filtering criteria outlined in [15]. Based on the number of non-zeros of the constraint matrices, we further categorize these 383 instances into three groups, as summarized in Table 1.
@@
Software. We implement cuPDLPx in both C and Julia. The Julia version utilizes CUDA.jl as the interface for working with NVIDIA CUDA GPUs, while the C version directly leverages low-level CUDA libraries for improved performance. We compare the performance of cuPDLPx against cuPDLP [1], HPR-LP v0.1.0 [3], Gurobi 12.0.4 [13], and cuPDLP-C v0.4.1 [2]. All Julia implementations and codes are timed after pre-compilation. For the Gurobi experiments, we set Method=6 to activate PDLP and GURO_PAR_PDHGGPU=1 to enable GPU acceleration.
Computing environment. We use an NVIDIA H100-SXM-80GB GPU, with CUDA 12.4. The experiments of cuPDLP, HPR-LP and cuPDLPx (Julia version) are performed in Julia 1.11.5.
Initialization. cuPDLPx uses all-zero vectors as the initial starting points.
Optimality termination criteria. cuPDLPx terminate when the relative KKT error is no greater than the termination tolerance $\epsilon\in(0,\infty)$:
@@
The termination criteria are checked for the original LP instance, not the preconditioned ones, so that the preconditioning does not impact the termination. In the experiment, we set $\epsilon=10^{-4}$ for moderate accuracy and $\epsilon=10^{-8}$ for high accuracy. For Gurobi, both absolute and relative feasibility tolerances are checked, and feasibility is considered satisfied once either tolerance is met. To align termination criteria across solvers, we fix the absolute feasibility tolerance to GURO_PAR_PDHGABSTOL=1e-9, while GURO_PAR_PDHGCONVTOL and GURO_PAR_PDHGRELTOL serve as the effective convergence criteria in Gurobi.
Time limit. We apply a time limit of 3600 seconds to small and medium-sized problems and 18000 seconds to large-scale problems for MIPLIB relaxation. A 15000-second time limit is used for Mittelmann’s benchmark.
@@
Tables 2 and 3 compare the performance of cuPDLP and cuPDLPx on 383 MIPLIB instances, with and without Gurobi presolve, respectively. Several key observations can be made:
  • Across all four settings, cuPDLPx consistently outperforms cuPDLP in terms of overall numerical performance, solving more instances and achieving notable reductions in running time.
  • The performance gain of cuPDLPx is especially significant for Small and Medium instances, while the improvement is less significant for Large instances, likely due to the high variance among the relatively small number of Large instances.
Table 4 reports the speedup of cuPDLPx over cuPDLP. In particular, the first row focuses on hard instances, defined as those requiring at least 10 seconds to solve using either cuPDLP or cuPDLPx . In various scenarios, cuPDLPx achieves a 2× to 5× speedup. The improvement is more substantial when Gurobi presolve is enabled compared to the case without presolve. The speedup is also more significant on harder instances. For example, in the high-accuracy setting with presolve, cuPDLPx achieves a 3.57× speedup on all instances and a 4.99× speedup on hard instances. These results demonstrate that the algorithmic enhancements of cuPDLPx are the primary driver of the observed 2×–5× improvements, while the C implementation yields an additional performance gain of about 20% compared to the Julia prototype.
$\epsilon$ Solver
@@
$\mathbf{10^{-4}}$ cuPDLP (Julia) 266 8.81 91 11.55 19 77.43 376 11.07
HPR-LP (Julia) 268 3.94 92 6.46 15 150.85 375 6.50
cuPDLPx (Julia) 269 3.50 93 5.51 19 62.04 381 5.24
cuPDLP-C 268 3.78 92 7.54 18 60.12 378 5.92
Gurobi-PDLP 267 4.74 92 7.67 17 103.47 376 7.14
cuPDLPx (C) 269 2.77 94 4.62 19 57.18 382 4.39
$\mathbf{10^{-8}}$ cuPDLP (Julia) 260 24.80 87 36.90 17 208.66 364 31.22
HPR-LP (Julia) 261 10.77 91 16.78 14 485.71 366 16.09
cuPDLPx (Julia) 264 8.94 91 16.43 17 194.91 372 13.27
cuPDLP-C 265 11.72 87 20.81 16 186.44 368 16.55
Gurobi-PDLP 261 12.22 89 19.77 17 388.87 367 17.76
cuPDLPx (C) 266 6.84 92 13.91 16 194.40 374 10.91
@@
$\mathbf{10^{-4}}$ cuPDLP (Julia) 268 5.11 92 9.03 19 26.74 379 6.75
HPR-LP (Julia) 268 2.21 94 3.88 19 15.32 381 3.08
cuPDLPx (Julia) 269 1.85 94 3.36 19 15.06 382 2.69
cuPDLP-C 268 1.99 94 4.25 19 13.25 381 2.95
Gurobi-PDLP 269 2.57 93 5.09 19 16.52 381 3.67
cuPDLPx (C) 269 1.41 94 2.75 19 14.07 382 2.19
$\mathbf{10^{-8}}$ cuPDLP (Julia) 264 18.50 90 29.40 19 63.68 373 22.42
HPR-LP (Julia) 268 5.85 92 11.90 19 45.63 379 8.32
cuPDLPx (Julia) 269 5.16 93 11.02 19 46.84 381 7.60
cuPDLP-C 264 9.28 89 16.05 18 46.70 371 11.96
Gurobi-PDLP 265 7.35 90 14.52 19 39.65 374 9.95
cuPDLPx (C) 269 3.98 92 9.60 19 43.09 380 6.28
@@
Instance Version Without Presolve With Presolve
$10^{-4}$ $10^{-8}$ $10^{-4}$ $10^{-8}$
Hard Julia 3.82 3.09 4.06 4.14
C 4.60 3.81 4.85 4.99
Overall Julia 2.11 2.35 2.51 2.95
C 2.52 2.86 3.08 3.57
Table 4: Summarization of cuPDLPx speedup (Julia and C versions) over cuPDLP. The “Hard” rows report speedup on instances where either solver takes more than 10 seconds, while the “Overall” rows summarize speedup across all instances.
Tables 5 and 6 summarize results on Mittelmann’s LP benchmark set with 49 instances. Consistent with the MIPLIB experiments, cuPDLPx achieves clear speedups over cuPDLP across all tolerances and presolve settings. At tolerance $10^{-4}$, cuPDLPx yields about a 3×–4× speedup without presolve, which increases to 6.8× with presolve. At the tighter tolerance $10^{-8}$, the improvements remain substantial, with about a 3× speedup without presolve and over 5× with presolve.
Solver $\mathbf{\epsilon=10^{-4}}$ $\mathbf{\epsilon=10^{-8}}$
Count Time Count Time
cuPDLP (Julia) 43 67.38 40 219.51
HPR-LP (Julia) 48 24.28 42 96.08
cuPDLPx (Julia) 48 22.78 44 78.62
cuPDLP-C 47 35.36 40 141.02
Gurobi-PDLP 44 32.47 39 153.94
cuPDLPx (C) 48 19.11 44 70.64
Table 5: Solve time in seconds and SGM10 on instances of Mittelmann’s LP benchmark set without presolve.
Solver $\mathbf{\epsilon=10^{-4}}$ $\mathbf{\epsilon=10^{-8}}$
Count Time Count Time
cuPDLP (Julia) 46 48.76 42 187.22
HPR-LP (Julia) 49 11.85 47 53.17
cuPDLPx (Julia) 49 9.30 48 38.96
cuPDLP-C 47 19.32 45 103.82
Gurobi-PDLP 46 21.95 43 93.29
cuPDLPx (C) 49 7.09 47 33.70
Table 6: Solve time in seconds and SGM10 on instances of Mittelmann’s LP benchmark set with Gurobi presolve.
5 Conclusions
In this paper, we present cuPDLPx , an enhanced GPU implementation of restarted Halpern PDHG for solving large-scale LPs. Numerical experiments on both MIPLIB relaxations and Mittelmann’s benchmark set show that cuPDLPx consistently outperforms cuPDLP, achieving 2×–6× speedups across different tolerances and presolve settings. In particular, the C implementation provides further performance improvements over the Julia prototype. These results highlight the effectiveness of algorithmic and implementation enhancements in exploiting GPUs for high-performance linear programming.
@@ Acknowledgement
Haihao Lu is partially supported by AFOSR Grant No. FA9550-24-1-0051 and ONR Grant No. N000142412735. Zedong Peng is partially supported by ONR Grant No. N000142412735. Jinwen Yang is partially supported by AFOSR Grant No. FA9550-24-1-0051.